OpenStudy (anonymous):

if f(x)=sqrt 36-x^2 then the range of value of f is given by?

OpenStudy (anonymous):

but how? @perl

OpenStudy (perl):

well you can graph it , thats one way. also you can reason it like this. sqrt(0) = 0, sqrt(1) = 1 , sqrt(4) = 2 , etc. and i realize that youre domain is -6 <= x < = 6 , so the range is not to infinity

OpenStudy (perl):

f(0) = sqrt( 36 -0^2) = sqrt(36) = 6 f(6) = 0 f(-6) = 0 so the range is [ 0 , 6 ]

OpenStudy (anonymous):

we have to put values of x?

OpenStudy (perl):

it helps, or you can graph it

OpenStudy (perl):

there is no simple way to find the range , unless you want to solve for x in terms of y, and then find the domain

OpenStudy (anonymous):

ok ty